OpenStudy (anonymous):

7(x − 1) = 5(x + 2) + 1

OpenStudy (anonymous):

1st step:expand and reduce both sides of the equation

OpenStudy (wwe123):

7(x − 1) = 5(x + 2) + 1 7x-7= 5x+10+1 2x=11+7 x=18/2 x=9